OpenStudy (anonymous):

A four-digit hexadecimal integer is written on a napkin such that the units digit is illegible. The first three digits are 4, A, and 7. If the integer is a multiple of 17_{10}, what is the units digit?

ganeshie8 (ganeshie8):

say the units digit is x : \[4A7x\] notice that 17 = 11 in hexadecimal, so the same divisibility rule for 11 works in hexadecimal too : \[(4 + 7) - (A+x)\] must be divisible by 17

OpenStudy (ikram002p):

Hi :)

ganeshie8 (ganeshie8):

Hello

ganeshie8 (ganeshie8):

\[(4A7x)_{16} \equiv 0 \pmod{17}\] \[4\times 16^3 + A\times 16^2 +7\times 16^1 + x\times 16^0 \equiv 0 \pmod{17}\] \(16\equiv -1 \pmod{17}\) so \[4\times (-1)^3 + A\times (-1)^2 +7\times (-1)^1 + x\times (-1)^0 \equiv 0 \pmod{17}\] \[-4+ A-7+x \equiv 0 \pmod{17}\]

OpenStudy (anonymous):

thanks a ton, gneshie8; so x= 1, this was easy.. as all ques. are after you solve them

ganeshie8 (ganeshie8):

yes x=1 :)